A chef is using a mixture of two brands of Italian dressing. The first brand contains 9% vinegar, and the second brand contains 14% vinegar. The chef wants to make 300 milliliters of a dressing that is 10% vinegar. How much of each brand should she use?
Answer by richwmiller(17219)

a+b=300,
0.09*a+0.14*b=0.1*300
a=300-b
0.09*(300-b)+0.14*b=30
27-0.09b+0.14*b=30
0.05*b=3
b=60
a=300-b
a=240
b=60
a=240 b=60